# Against The Achilles’ Heel: A Survey on Red Teaming for Generative Models

### v1

@@ 4.1.4 Model Manipulation
Till now, we have introduced attack methods created by exploiting LLM’s capabilities under the assumption that we cannot manipulate LLMs themselves. However, as access to LLM services expands and more open-source versions become available, in this section, we turn our attention to methods that involve adjusting the hyper-parameters or weights of LLMs to mount attacks.
@@
Languages model outputs vary tremendously under different decoding methods and hyper-parameters. Inie et al. (2023) noted that changing temperature increases the randomness of tokens allowing more attack possibilities. Huang et al. (2023b) explored this phenomenon in depth and highlighted the potential risks. They noticed that many open-source LLMs only perform alignment under default decoding sampling methods and hyper-parameters. By sampling temperatures, changing the parameters $K$ and $p$ in Top-$K$ and Top-$p$ samplings respectively, they demonstrated that the attack success rate increased significantly from 9% to more than 95% and decoding with penalty and constraints further elevates the ASR metric.
In a notable study, researchers demonstrated how adjusting the model’s probability to favor affirmative responses, such as “Sure, here is”, can make the model agree to dangerous instructions. This method was effectively demonstrated in the study (Zhang et al., 2023b). Fort (2023) further discusses the “scaling law” between the length of manipulated tokens and their maximum controllable length of subsequently generated content. Further discussion about this affirmative suffix strategy is shown in Sec. 4.1.1.
@@
A key strategy in this domain is the use of interference vectors. When introduced during the model’s inference phase, these interference vectors can significantly alter its output and steer it toward a specific direction effectively. Wang & Shu (2023) demonstrates how an extra layer can interfere with and redirect the model’s inference pathway. Similarly, Li et al. (2024d) explores the impact of embedding interference vectors directly into the inference process. Both approaches underline the vulnerability of LLMs to manipulation through their inference mechanisms.
With the rise of automatic prompt optimization (Yang et al., 2023a; Pryzant et al., 2023; Zhou et al., 2022), some studies have utilized this method to modify malicious prompts, enabling them to bypass model safety constraints and trigger malicious outputs (Chao et al., 2023; Mehrotra et al., 2023).
@@
With the growth of the open source community and the widespread adoption of open-source LLMs, fine-tuning has become a prevalent method for tailoring services to specific needs. The official documentation for the open-sourced Llama2 recommends fine-tuning for developing custom products, with the aim of enhancing the model’s performance for particular applications (Peng et al., 2023). Likewise, OpenAI has introduced APIs for fine-tuning GPT3.5 with custom datasets. This move highlights findings from their private beta, indicating that fine-tuning has enabled users to significantly enhance the model’s effectiveness across various applications (Meta, 2023b).
@@
However, plenty of recent research has demonstrated that the safety alignment of LLMs can be compromised by fine-tuning with only a few training examples (Yang et al., 2023b; Qi et al., 2023b; Chen et al., 2023b; Lermen et al., 2023; Gade et al., 2023; Chen et al., 2023b; Zhan et al., 2023). These studies typically involve collecting a small dataset of harmful instructions and responses, fine-tuning safety-aligned LLMs with these datasets, and then evaluating the models’ performance for both their helpfulness and harmfulness. Such experiments have been conducted on several open-source models including Falcon, Baichuan, IntermLM, Llama and the widely recognized safest open-source model, Llama2, as well as the most powerful models currently accessible only via API: GPT3.5 and GPT4. Despite employing various general-purpose and safety evaluation benchmarks, the results show a consistent trend: the performance on general benchmarks does not diminish after fine-tuning, whereas the rate of harmfulness increases significantly, from less than 10% to over 80%.
@@
In addition to identifying fine-tuning as a method to bypass the safeguards of LLMs, researchers have explored various aspects of fine-tuning that impact the success rate of breaching these safeguards. Notably, several works underscore the low cost of jailbreak via fine-tuning, achieved by using a small amount of instruction-tuning data (Zhan et al., 2023; Yang et al., 2023b) and parameter-efficient fine-tuning techniques (Lermen et al., 2023). Qi et al. (2023b) indicate that higher learning rates and smaller batch sizes generally result in greater degradation of safety and increased harmfulness rates. This phenomenon may be attributed to large and unstable gradient updates that cause significant deviations in safety alignment. Additionally, Yang et al. (2023b) demonstrates that fine-tuning with single-turn English data can lead to the degradation of the model’s safeguards when transitioning to multiple-turn dialogue in other languages.
@@
Fine-tuning with explicitly harmful data can breach the safety limits of a model. However, fine-tuning with data that is not overly harmful can also potentially compromise the model’s safety alignment. This issue has been the subject of extensive discussion in various studies. Zhan et al. (2023) reveal that employing multi-turn, non-factual, in-context learning examples can prompt the model to generate harmful outputs. Specifically, their method compels the model to accept statements like “1+1 is 3” and “the earth is flat”. Yang et al. (2023b) explores three levels of data: (1) explicitly harmful data, (2) implicitly harmful data, and (3) benign data. For the first two, the authors observe a noticeable deterioration in safety. Furthermore, Yang et al. (2023b) demonstrates that even using benign datasets (level 3), without malicious intent, mere fine-tuning with such datasets, which are purely utility-oriented, could compromise the safety alignment of LLMs. They conducted experiments on widely used instruction-tuning datasets like Alpaca (Taori et al., 2023) and Dolly (Conover et al., 2023), as well as the vision-language model using LLaVA-instruct (Liu et al., 2023b), and consistently observed a decline in safety across all evaluated cases post fine-tuning. The authors attribute this to the delicate balance required between the safety/harmlessness and capability/helpfulness of LLMs. Reckless fine-tuning could upset this balance, such as fine-tuning an aligned LLM on a utility-oriented dataset, which might divert models from their objective of harmlessness. Moreover, there’s a risk of catastrophic forgetting of the model’s initial safety alignment during fine-tuning, as highlighted by prior research (Kirkpatrick et al., 2017; Luo et al., 2023).
@@
Apart from those studies mentioned above, which work on general safety problems, Chen et al. (2023b) focused on unintended Personal Identifiable Information (PII) disclosure of LLMs after fine-tuning. They reveal a startling propensity for models, exemplified by GPT3.5, to transition from a state of PII non-disclosure to one where they unveil a significant volume of hitherto safeguarded PII, with only minimal fine-tuning intervention.
@@ 4.2 Attack Searchers
  • State Space is the set containing all possible states. In DFS, the state space is all nodes in the graph. In red teaming searchers, we identify two types of state spaces: prompts and suffixes. Prompt searchers find prompts that induce dangerous content, while those for suffixes intend to find a transferable suffix to various prompts, triggering their jailbreak behavior.
  • Search Goal is the goal of the searcher. In DFS, the search goal can be finding a node with certain attributes. In red teaming searchers, the ultimate goal is to jailbreak language models with attack prompts. These searchers are often implemented as a classifier or string matching Sec. 5.3. Some work employs proxy goals to enable certain kinds of search operations or to utilize existing prompts. These proxy goals often stem from strategies discussed above, e.g. completion compliance and instruction indirection.
  • Search Operation is how the searcher iterates on the current state and approaches its search goal. In DFS, the search operation is traversing to the next node. In red teaming searchers, search operations abound because of the flexibility of the state space. Common operations include language model rewrites, greedy coordinate gradient (GCG) (Zou et al., 2023; Zhu et al., 2023), genetic algorithm (Lapid et al., 2023; Liu et al., 2023d), and Reinforcement Learning.
@@
We list all searchers in Table 1. Decoupling elements from searchers reveals a large design space for future exploration. For example, The AutoDAN approach uses GCG (Sec. 4.2.3) as a search operation combined with the proxy goal of affirmative suffixes; the goal can be changed to creating context-switching content, replying with ciphers, personifications or other possible strategies discussed above.
4.2.1 State Space
@@
Puzzler (Chang et al., 2024) Prompted LLM Composition
GUARD (Jin et al., 2024) Prompted LLM Composition + LLM Mutation
PAIR (Chao et al., 2023) Prompted LLM LLM Mutation
TAP (Mehrotra et al., 2023) Prompted LLM LLM Mutation
GBRT (Wichers et al., 2024) Safety Classifier Gumbel Sampling
Explore, Establish, Exploit (Casper et al., 2023) Roberta RL(PPO)
@@
JADE (Zhang et al., 2023c) Instruction Indirection Grammar Tree Mutation
AutoDAN-GA (Liu et al., 2023d) Instruction Indirection Genetic Algorithm + LLM Mutation
Prompt Packer (Jiang et al., 2023d) Instruction Indirection Composition
FuzzLLM (Yao et al., 2023a) Instruction Indirection Composition + LLM Mutation
HouYI (Liu et al., 2023g) Instruction Indirection Composition
GPTFuzzer (Yu et al., 2023a) Context Switching LLM Mutation
SimpleRephrase (Takemoto, 2024a) Instruction Indirection LLM Mutation
Embedding Attack (Schwinn et al., 2023) Affirmativeness Optimization
CPAD (Liu et al., 2023a) Instruction Indirection LLM Mutation
COLD (Guo et al., 2024) Affirmative Suffixes COLD Decoding
@@
PAL (Sitawarin et al., 2024) Affirmative Suffixes GCG + Filtering
AutoDAN-I (Zhu et al., 2023) Affirmative Suffixes + Readability GCG
SESAME (Lapid et al., 2023) Instruction Indirection Genetic Algorithm
TrojLLM (Xue et al., 2023) RL Reward Function RL
Table 1: We list papers whose method can be framed as a searching problem here.
@@
One of the attributes of prompt searchers is the use of templates: existing red teaming prompts discovered in the wild. They compose or mutate existing templates before inserting attack topics to create the final prompt. Filtering is needed as some of the prompts may not be effective, after which successful prompts will often be added back to the template library for future reference. Examples include CPAD (Liu et al., 2023a), the HouYi framework (Liu et al., 2023g), and AutoDAN-GA (Liu et al., 2023d). [^1] We found that multiple papers proposed algorithms named AutoDAN, so we added suffixes to distinguish them. CPAD builds 5 strategies to increase diversity from replacing keywords to adding role play, whereas HouYi divides the prompt into modules and generates them separately. These modules include the main framework, disruptors for context switching (see Sec. 4.1.1) and separators for Instruction Indirection (see Sec. 4.1.2).
Prompt searchers without templates directly work with an attacking prompt. As the initial prompt contains direct instructions that are easily discerned, these searchers mutate these prompts with their search operations to add more complexity in hopes that the resulting query successfully attacks the target models. Searchers like Pair (Chao et al., 2023) and TAP (Mehrotra et al., 2023) falls into this category. Both of these searchers employ language models to search through the state space by mutating them to become more indirect. TAP is more involved in their search process by using Tree-of-Thoughts (Yao et al., 2023b) that enables more exploration. JADE (Zhang et al., 2023c) builds parsing trees of the query and applies grammar tree mutation to add more complexity.
@@
Searchers of suffixes lean on prompt optimization by operating on a suffix instead of the whole prompt to achieve maximum generalizability, which has proven effective in red teaming language models. The key challenge lies in finding the optimization goal while handling the exponential complexity of suffix searching. Typical suffix searchers are AutoDAN-I (Zhu et al., 2023), PAL (Sitawarin et al., 2024) and TrojLLM (Xue et al., 2023), employing various proxy goals and optimization algorithms to combat complexity.

### v2

@@ 4.1.4 Model Manipulation
To this point, we have introduced attack methods created by exploiting LLM’s capabilities under the assumption that we cannot manipulate the LLMs themselves. However, as access to LLM services expands and more open-weight and fully open-source Liu et al. (2023j) models become available, attack methods that involve adjusting the hyper-parameters or weights of LLMs are becoming more prevalent. While we consider the differentiation between open-weight and open-source models to be critical for many applications (for truly open-source models, all data, code, logs, and checkpoints are made available, whereas with open-weight models, only the learned parameters are released), for the purposes of this section we refer to them collectively as “open models”.
@@
Language model outputs vary tremendously under different decoding methods and hyper-parameters. Inie et al. (2023) noted that changing the temperature increases the randomness of generated tokens, allowing more attack possibilities. Huang et al. (2023c) explored this phenomenon in depth and highlighted the potential risks. They noticed that many open LLMs only perform alignment under default decoding sampling methods and hyper-parameters. By sampling temperatures, changing the parameters $K$ and $p$ in Top-$K$ and Top-$p$ samplings respectively, they demonstrated that the attack success rate increased significantly from 9% to more than 95% and decoding with penalty and constraints further elevates the ASR metric.
In a notable study, researchers demonstrated how adjusting the model’s probability to favor affirmative responses, such as “Sure, here is”, can make the model agree to dangerous instructions. This method was effectively demonstrated in the study Zhang et al. (2023b). Fort (2023) further discusses the “scaling law” between the length of manipulated tokens and their maximum controllable length of subsequently generated content. Further discussion about this affirmative suffix strategy is shown in Section 4.1.1.
@@
A key strategy in this domain is the use of interference vectors. When introduced during the model’s inference phase, these interference vectors can significantly alter its output and steer it in a specific direction. Wang & Shu (2023) demonstrates how an extra layer can interfere with and redirect the model’s inference pathway. Similarly, Li et al. (2024f) explores the impact of embedding interference vectors directly into the inference process. Both approaches underline the vulnerability of LLMs to manipulation through their inference mechanism.
With the rise of automatic prompt optimization Yang et al. (2023a); Pryzant et al. (2023); Zhou et al. (2023), some studies have utilized this method to modify malicious prompts, enabling them to bypass model safety constraints and trigger malicious outputs Chao et al. (2023); Mehrotra et al. (2023).
@@
With the increasing number of open models, fine-tuning has become a prevalent method for tailoring services to specific needs. The official documentation for the open-weight Llama2 recommends fine-tuning for developing custom products, with the aim of enhancing the model’s performance for particular applications Peng et al. (2023). Likewise, OpenAI has introduced APIs for fine-tuning the closed-weight GPT3.5 with custom datasets. This move highlights findings from their private beta, indicating that fine-tuning has enabled users to significantly enhance the model’s effectiveness across various applications Meta (2023b).
@@
However, there is a growing body of recent research that has demonstrated that the safety alignment of LLMs can be compromised by fine-tuning with only a few training examples Yang et al. (2023b); Qi et al. (2023); Chen et al. (2023b); Lermen et al. (2023); Gade et al. (2023); Chen et al. (2023b); Zhan et al. (2023). These studies typically involve collecting a small dataset of harmful instructions and responses, fine-tuning safety-aligned LLMs with these datasets, and then evaluating the models’ performance for both helpfulness and harmfulness. Such experiments have been conducted on several open models including Falcon, Baichuan, IntermLM, Llama, and Llama2, as well as models currently accessible only via API: GPT3.5 and GPT4. Despite employing various general-purpose and safety evaluation benchmarks, the results show a consistent trend: the performance on general benchmarks does not diminish after fine-tuning, whereas the rate of harmfulness increases significantly, from less than 10% to over 80%.
@@
In addition to identifying fine-tuning as a method for bypassing the safeguards of LLMs, researchers have explored various aspects of fine-tuning that impacts on the success rate of breaching these safeguards. Notably, much work has underscored the low cost of jailbreak via fine-tuning, achieved by using a small amount of instruction-tuning data Zhan et al. (2023); Yang et al. (2023b) and parameter-efficient fine-tuning techniques Lermen et al. (2023). Qi et al. (2023) indicate that higher learning rates and smaller batch sizes generally result in greater degradation of safety and increased harmfulness rates. This phenomenon may be attributed to large and unstable gradient updates that lead to significant deviations in safety alignment. Additionally, Yang et al. (2023b) demonstrate that fine-tuning with single-turn English data can lead to the degradation of the model’s safeguards when transitioning to multiple-turn dialogue in other languages.
Attack Category Sub-category Key Methods/Techniques
Completion
Compliance
Affirmative Suffixes Using phrases like ”Sure, here is” or ”Hello” Wei et al. (2023a)
Long suffixes mimicking assistant responses Rao et al. (2023)
Response inclination analysis Du et al. (2023b)
[…]
Semantic separators and HouYi framework Liu et al. (2023g)
Task switching techniques Inie et al. (2023)
In-context Learning Chain of utterances Bhardwaj & Poria (2023b)
In-context attacks Wei et al. (2023b)
Contextual interaction attacks Cheng et al. (2024)
Instruction
Indirection
Input Euphemisms Veiled expressions Xu et al. (2023d)
Socratic questioning Inie et al. (2023)
Altered sentence structures Ding et al. (2023)
Output Constraints Style constraints (Wikipedia, JSON) Wei et al. (2023a)
Task Constraints & safety behaviors Fu et al. (2023b)
Refusal suppression Schulhoff et al. (2023)
Virtual Simulation DeepInception scenario simulation Li et al. (2023e)
Program execution simulation Liu et al. (2023h)
Payload splitting Kang et al. (2023)
Generalization
Glide
Languages Multilingual attack strategies Wang et al. (2023c)
Low-resource language exploitation Deng et al. (2023d)
Cross-lingual safety analysis Shen et al. (2024b)
Ciphers Word substitution (ROT13, Caesar) Yuan et al. (2023b)
ASCII art encoding Jiang et al. (2024a)
SelfCipher & auto-obfuscation Wei et al. (2023a)
Personification Role play & persona modulation Shah et al. (2023)
Psychological manipulation Zeng et al. (2024a)
Privilege escalation Liu et al. (2023h)
Model
Manipulation
Decoding Manipulation Temperature & sampling manipulation Huang et al. (2023c)
Probability control Zhang et al. (2023b)
Weak-to-strong transfer Zhao et al. (2024d)
Activations Manipulation Interference vectors Wang & Shu (2023)
Embedding manipulation Li et al. (2024f)
Automatic prompt optimization Chao et al. (2023)
Model Fine-tuning Small dataset fine-tuning Yang et al. (2023b)
Parameter-efficient tuning Lermen et al. (2023)
PII disclosure risks Chen et al. (2023b)
Table 1: Summary of jailbreak approaches organized by attack categories.
Fine-tuning with explicitly harmful data can breach the safety limits of a model. However, fine-tuning with data that is not overly harmful can also potentially compromise the model’s safety alignment. This issue has been the subject of extensive discussion in various studies. Zhan et al. (2023) reveal that employing multi-turn, non-factual, in-context learning examples can prompt the model to generate harmful outputs. Specifically, their method compels the model to accept statements like “1+1 is 3” and “the earth is flat”.
@@
Yang et al. (2023b) explore three levels of data: (1) explicitly harmful data, (2) implicitly harmful data, and (3) benign data. For the first two, the authors observe a noticeable deterioration in safety. They show that merely fine-tuning with purely utility-oriented benign datasets without malicious intent (level 3), could compromise the safety alignment of LLMs. They conducted experiments on widely used instruction-tuning datasets like Alpaca Taori et al. (2023) and Dolly Conover et al. (2023), as well as a vision-language model using LLaVA-instruct Liu et al. (2023b), and consistently observed a decline in safety across all evaluated cases post fine-tuning. The authors attribute this to the delicate balance between safety/harmlessness and capability/helpfulness. Reckless fine-tuning could upset this balance, such as fine-tuning an aligned LLM on a utility-oriented dataset, which might divert models from their objective of harmlessness. Moreover, there’s a risk of catastrophic forgetting of the model’s initial safety alignment during fine-tuning, as highlighted by prior research Kirkpatrick et al. (2016); Luo et al. (2023).
@@
Apart from the studies mentioned above, which work on general safety problems, Chen et al. (2023b) focused on unintended Personal Identifiable Information (PII) disclosure of LLMs after fine-tuning. They reveal a startling propensity for models, exemplified by GPT3.5, to transition from a state of PII non-disclosure to one where they unveil a significant volume of hitherto safeguarded PII, with only minimal fine-tuning intervention.
Section Summary
This section outlined various attack strategies that exploit model capabilities, delving into methods that leverage autoregressive generation, instruction-tuned abilities, and domain generalization to bypass safety measures. The strategies use specific model capabilities like completion compliance, where attackers exploit the model’s tendency to complete prompts, and context switching, where distractions are introduced to override safety guidelines. Instruction-based attacks manipulate model responses through euphemisms or format constraints, while generalization allows exploitation across languages and encoded formats. Additionally, model manipulation techniques, including decoding manipulation and fine-tuning, demonstrate how model parameters or weights can be adjusted to amplify vulnerabilities. These methods collectively reveal potential risks within language models, particularly as they scale and are deployed across diverse applications. A summary of jailbreak approaches organized by attack categories is listed in Table 1.
@@ 4.2 Attack Searchers
  • State Space is the set containing all possible states. In DFS, the state space is all nodes in the graph. In red teaming searchers, we identify two types of state spaces: prompts and suffixes. Prompt searchers find prompts that induce dangerous content, while those for suffixes intend to find a suffix that transfers to various prompts, triggering their jailbreak behavior.
  • Search Goal is the goal of the searcher. In DFS, the search goal can be finding a node with certain attributes. In red teaming searchers, the ultimate goal is to jailbreak language models with attack prompts. These searchers are often implemented as a classifier or string matching Section 6.3. Some work employs proxy goals to enable certain kinds of search operations or to use existing prompts. These proxy goals often stem from strategies discussed above, e.g., completion compliance and instruction indirection.
  • Search Operation is how the searcher iterates on the current state and approaches its search goal. In DFS, the search operation is to traverse to the next node. In red teaming searchers, search operations abound because of the flexibility of the state space. Common operations include language model rewrites, greedy coordinate gradient (GCG) Zou et al. (2023b); Zhu et al. (2023), genetic algorithm Lapid et al. (2023); Liu et al. (2023d), and reinforcement learning.
We list all searchers in Table 2. Decoupling elements from searchers gives rise to a large design space for future exploration. For example, the AutoDAN approach uses GCG (Section 4.2.3) as a search operation combined with the proxy goal of affirmative suffixes; the goal can be changed to creating context-switching content, replying with ciphers, personifications, or other possible strategies discussed above.
@@
Puzzler Chang et al. (2024) Prompted LLM Composition
GUARD Jin et al. (2024) Prompted LLM Composition + LLM Mutation
PAIR Chao et al. (2023) Prompted LLM LLM Mutation
TAP Mehrotra et al. (2023) Prompted LLM LLM Mutation
GBRT Wichers et al. (2024) Safety Classifier Gumbel Sampling
EEE Casper et al. (2023) Roberta RL (PPO)
@@
JADE Zhang et al. (2023c) Instruction Indirection Grammar Tree Mutation
AutoDAN-GA Liu et al. (2023d) Instruction Indirection Genetic Algorithm + LLM Mutation
Prompt Packer Jiang et al. (2023d) Instruction Indirection Composition
FuzzLLM Yao et al. (2023a) Instruction Indirection Composition + LLM Mutation
HouYI Liu et al. (2023g) Instruction Indirection Composition
GPTFuzzer Yu et al. (2023a) Context Switching LLM Mutation
SimpleRephrase Takemoto (2024a) Instruction Indirection LLM Mutation
Embedding Attack Schwinn et al. (2023) Affirmativeness Optimization
CPAD Liu et al. (2023a) Instruction Indirection LLM Mutation
COLD Guo et al. (2024) Affirmative Suffixes COLD Decoding
@@
PAL Sitawarin et al. (2024) Affirmative Suffixes GCG + Filtering
AutoDAN-I Zhu et al. (2023) Affirmative Suffixes + Readability GCG
SESAME Lapid et al. (2023) Instruction Indirection Genetic Algorithm
TrojLLM Xue et al. (2023) RL Reward Function RL
Table 2: List of methods that can be framed as searching problems.
4.2.1 State Space
@@
One attribute of prompt searchers is the use of templates: existing red teaming prompts discovered in the wild. They compose or mutate existing templates before inserting attack topics to create the final prompt. Filtering is needed as some of the prompts may not be effective, after which successful prompts will often be added back to the template library for future reference. Examples include CPAD Liu et al. (2023a), the HouYi framework Liu et al. (2023g), and AutoDAN-GA Liu et al. (2023d). [^2] We found that multiple papers proposed algorithms named AutoDAN, so we added suffixes to distinguish them. CPAD builds 5 strategies to increase diversity from replacing keywords to adding role play, whereas HouYi divides the prompt into modules and generates them separately. These modules include the main framework, disruptors for context switching (see Section 4.1.1) and separators for Instruction Indirection (see Section 4.1.2).
Prompt searchers without templates work directly with an attacking prompt. As the initial prompt contains direct instructions that are easily discerned, these searchers mutate these prompts with their search operations to add more complexity in the hopes that the resulting query successfully attacks the target models. Searchers like Pair Chao et al. (2023) and TAP Mehrotra et al. (2023) fall into this category. Both employ language models to search through the state space by mutating them to become more indirect. TAP is more involved in their search process by using Tree-of-Thoughts Yao et al. (2023b), enabling more exploration. JADE Zhang et al. (2023c) builds parse trees of the query and applies mutation of the syntactic tree to add more complexity.
@@
Searchers of suffixes lean on prompt optimization by operating on a suffix instead of the whole prompt to achieve maximum generalizability, which has proven effective in red teaming language models. The key challenge lies in finding the optimization goal while handling the exponential complexity of suffix searching. Typical suffix searchers include GCG Zou et al. (2023b), AutoDAN-I Zhu et al. (2023), PAL Sitawarin et al. (2024) and TrojLLM Xue et al. (2023), employing various proxy goals and optimization algorithms to combat complexity.